# middlelinecolor=black, middlelinewidth=0.5pt, backgroundcolor=gray!10, innerleftmargin=4pt, innerrightmargin=4pt, innertopmargin=4pt, innerbottommargin=4pt, roundcorner=4pt

### v1

@@ 4.1 Results
DML (Zhang et al., 2018) 74.83 70.55 70.24 72.86 66.30 74.52
ONE (Zhu et al., 2018) 74.68 70.77 70.43 72.01 66.26 -
SHAKE (Li & Jin, 2022) 75.78 - 71.62 73.85 68.81 76.42
TAKD (Mirzadeh et al., 2019) 75.04 - 70.77 73.67 - -
AEKD (Du et al., 2020) 75.68 71.36 71.25 74.75 68.39 76.34
EBKD (Kwon et al., 2020) 74.10 68.24 76.61
DGKD (Son et al., 2021) 76.24 - 71.92 74.40 - -
CA-MKD (Zhang et al., 2022) - - - 74.30 69.41 77.94
@@
T: RG-Y320, 4 Peers
Student NOKD AVER Ours
R32 73.31 74.60 75.20
(c) Results on Pose Estimation with four mentors. We report PCKh for MPII and AP for COCO.
@@
Teacher HRNet-W32-D RTMP-L*
@@
CIFAR-100 Classification with Multiple Mentors. We compare our approach with online and offline strategies involving multiple mentors in Table 2(a). In addition to the teacher named in the table, our classroom uses five additional peers, as defined in Appendix A. ClassroomKD performs better than online methods like DML, ONE, and SHAKE, as well as the baseline offline method (AVER) in almost all cases. Our method also outperforms methods like TAKD, AEKD, and DGKD, which are specifically designed to address the capacity gap and error accumulation issues. In particular, the AEKD method achieves significantly poor results despite using four more mentors than us. We also have a comparable performance with the state-of-the-art CA-MKD method for networks they report.
ImageNet Classification with Multiple Mentors: In the ImageNet experiments, ClassroomKD demonstrates its scalability and effectiveness on a larger dataset. As shown in Table 2(c), ClassroomKD outperforms traditional KD and other multi-teacher methods, achieving higher top-1 and top-5 accuracy scores. This indicates that ClassroomKD maintains its superiority even as the complexity and size of the dataset increase, underscoring its robustness and adaptability.
Pose Estimation with Multiple Mentors. We also assess ClassroomKD’s performance on 2D human pose estimation tasks using the COCO Keypoints and MPII Human Pose datasets. Table 2(c) compares ClassroomKD against a simple multi-teacher baseline regarding keypoint detection accuracy and overall pose estimation performance. ClassroomKD achieves higher Average Precision (AP) scores, demonstrating its ability to effectively transfer structured knowledge from multiple mentors to the student model. This highlights ClassroomKD’s versatility and effectiveness beyond image classification tasks, extending its applicability to complex, structured prediction problems.
4.2 Ablation Studies
@@
Table 3: Ablation study to assess the contribution of different components of ClassroomKD.
(a) Role of Multiple Mentors. Single-teacher distillation slightly improves student performance compared to vanilla training. Adding intermediate mentors (peers) and using adaptive distillation further enhances learning.
Student Teacher Peers Adaptive Distillation Top-1 Accuracy
@@
4.2.1 Classroom Size and Composition
@@
11112222333344445555666664646464666666666868686863.3564.7164.7266.5767.0867.53Number of MentorsStudent Top-1S: EN-B0, T: EN-B0, 5 Same Peers
@@
Figure 4: Effect of Classroom Size and Composition. We investigate how the number of mentors, their architectures, and performance differences affect learning.
@@
(a) Diversity in mentor architectures. Using diverse mentor architectures improves distillation performance compared to a homogeneous setup, even when the total parameter count of the diverse mentors is lower. This indicates that architectural diversity provides valuable learning signals.
@@
4.2.2 Temperature in Mentoring Module
We explore the role of adaptive temperature ( $\tau$) in the Mentoring Module and its impact on bridging the capacity gap between classroom networks. Our approach adjusts the temperature dynamically based on the student’s learning progress, with higher $\tau$ values at the start to accommodate the larger capacity gap, which gradually decreases as the student’s understanding improves. This adaptive strategy allows mentors to effectively “slow down” the teaching process during early stages and accelerate it later, ensuring effective knowledge transfer.
@@ 4.2.2 Temperature in Mentoring Module
Table 5: Temperature adaption strategy. We compare our temperature adaptation method to DTKD (Wei & Bai, 2024) by replacing our mentoring module with their dynamic temperature computation. Our approach outperforms DTKD with $\tau=12$ and remains comparable or better even when DTKD’s base temperature is separately tuned to $\tau=8$.
@@
DTKD DTKD - 69.10 72.05
Ours DTKD 8 69.80 72.47
Ours DTKD 12 68.87 72.36
Ours Ours 12 69.84 72.65
@@
Comparison with DTKD. We further evaluate our approach against DTKD’s dynamic temperature strategy by incorporating their method into our mentoring module. While DTKD shows minor improvements with $\tau=8$ for MobileNetV2, it falls short when using $\tau=12$ or for other student models such as ResNet20. This highlights DTKD’s dependence on the initial $\tau$ value and its inability to fully address the capacity gap between the teacher and student. In contrast, our method adapts more effectively across different student and mentor configurations, achieving consistently better results by tailoring the teaching strategy dynamically throughout training.
4.2.3 Ranking Strategies in KF Module
@@
Table 6: Choice of Ranking Strategy. We compare three ranking methods. Here, we employ different networks for peers. (A) we use class probabilities as $\alpha$, $\beta$. (B) we employ class ranks as $\alpha$, $\beta$. (C) We use a dynamically calculated $\lambda$. We observe that using ranks as loss weights improves student network performance compared to probabilities.
Teacher WR40x2 R110 R110 R56 VGG13
Student WR16x2 R20 R32 R20 VGG8
Method B 75.42 71.94 74.28 72.56 73.58
Ours 76.51 72.45 74.60 72.65 74.51
Teacher VGG13 R32×4 R32×4 R50 WR40x2
Student MBV2 SN-V2 SN-V1 MBV2 SN-V1
Method B 68.52 75.71 75.08 69.78 75.96
Ours 69.84 76.52 74.84 70.15 75.05
@@ 5 Conclusion
We presented ClassroomKD, a novel knowledge distillation framework that mimics a classroom environment, where a student learns from a diverse set of mentors. By selectively integrating feedback through the Knowledge Filtering (KF) Module and dynamically adjusting teaching strategies with the Mentoring Module, ClassroomKD ensures effective knowledge transfer and mitigates the issues of error accumulation and capacity gap. Our approach significantly improves the student’s performance in classification and pose estimation tasks, consistently outperforming traditional distillation methods. Limitations. While we demonstrated the efficacy of ClassroomKD on image classification and human pose estimation, its application to other domains and more complex tasks, such as object detection and segmentation, presents a promising avenue for future work. Despite the improvements, the framework introduces complexity, especially with respect to the mentor ranking and teaching adjustments, which can require careful tuning. Future work will explore further optimizations and expand the framework’s utility to broader tasks.
Appendix A Training Protocols
@@
Table 7: Mentor Configurations.We show the set of models used in all our experiments along with their respective top-1 accuracies and the ensemble performance. The size of the mentors, if all the peers were replaced by the teacher ($(n+1))t$), the size of the current mentors ($1tnp$), and student size are also mentioned. Model abbreviations: MB: MobileNet, SN: ShuffleNet, R: Resnet, W: WRN, EN: EfficientNet, SQ: SqueezeNet, RP: RTMPose, HR: HRNet, LHR: LiteHRNet, RG: RegNet
@@
Hardware and Software Configuration. We trained most of our CIFAR-100 experiments on a single V100-16GB GPU. The time required for an experiment ranged between 4 and 4.5 hours on average. We build our code on top of Image Classification SOTA repository [^1] https://github.com/hunto/image_classification_sota/ and MMPose, and use pretrained models from these libraries as our mentors.
Appendix B Future Direction: ClassroomKD and Dataset Distillation
@@
Dataset distillation aims to create small, synthetic datasets that enable neural networks to achieve comparable performance to those trained on the original, much larger datasets. This approach reduces computational costs and storage requirements while maintaining model generalization. By optimizing a small set of representative training samples, a distilled dataset $S$ is generated such that a model trained on $S$ performs well on the original dataset $\mathcal{T}$. In our experiments, we use FRePo (Zhou et al., 2022) to create a distilled CIFAR-100 dataset, reducing each class to only 10 samples (Figure 7). Of these, 7 images per class are used for training, while the remaining 3 are used for testing.
@@
Appendix C Analysis and Additional Results
C.1 Per-Class Performance Improvement
Figure 8: Comparison of per-class performance gain over the NOKD baseline. With ClassroomKD (left), the distilled model improves performance on 86 classes. With multi-teacher KD without mentor ranking (right), significantly fewer classes improve, the absolute improvement is smaller, and the remaining classes experience larger performance degradation (red bars). This highlights the impact of our dynamic strategies in improving performance across different classes.
C.2 Softmax for Ranking Mentor
Figure 9: Softmax for Ranking Mentors. We plot the softmax probabilities (y-axis) against the class indices (x-axis) of all models in a 1t5p classroom used. On the left, we show the distribution of confidence in model predictions. We zoom in on the target class on the right. The mentors with less softmax probability (or confidence) than the student (blue line) are deactivated to reduce error accumulation observed in traditional multi-teacher distillation methods.
C.3 Pose Estimation Qualitative Results
(a) NOKD
(b) AVER
(c) Ours
Figure 10: Qualitative comparison of LiteHRNet-18 student trained (a) without distillation, (b) with six teachers in a naive manner, and (c) with six teachers using ClassroomKD.
Appendix D ClassroomKD Algorithm
Algorithm 1 ClassroomKD
0:  Input batch $\mathbf{x}$
0:  Ground truth labels $\mathbf{y}$
0:  Student $s$
0:  Mentors $\mathbb{M}\leftarrow\{t\}\cup\{p_{i}\}_{i=1}^{n}$
0:  $\beta$: weight of distillation loss
3:  $\mathcal{L}\leftarrow 0$ // Initialize total loss
4:  mask $\leftarrow$ onehot($\mathbf{y}$)
5:  $\mathbb{C}\leftarrow\{s\}\cup\mathbb{M}$
6:  for $m\in\mathbb{C}$ do
7:     $\hat{\bm{y}}^{m}\leftarrow m(\bm{x})$ // Get predictions from model $m$
8:     $\bm{p}^{m}\leftarrow\mathrm{softmax}(\hat{\bm{y}}^{m})$ // Convert logits to probabilities
9:     $\bm{p}^{m}_{\text{gt}}\leftarrow\bm{p}^{m}\cdot\text{mask}$ // Isolate probabilities assigned to ground truth
[…]
11:     $\text{weights}[m]\leftarrow w^{m}$ // Store weight for model $m$
12:  end for
[…]
16:  for $m\in\mathbb{M}$ do
17:     if $\text{ranks}[m]>\text{ranks}[s]$ then
18:        $\tau^{m}\leftarrow\frac{\text{ranks}[m]-\text{ranks}[s]}{\text{ranks}[m]}$
19:        $\mathcal{L}_{\mathrm{distill}}\leftarrow\mathrm{KL}(\hat{\bm{y}}^{m},\hat{\bm{y}}^{s},\tau^{m})$
20:        $\mathcal{L}_{\mathrm{distill}}\leftarrow\text{ranks}[m]\cdot\mathcal{L}_{\mathrm{distill}}$
21:     else
22:        $\mathcal{L}_{\mathrm{distill}}\leftarrow 0$
23:     end if
24:     $\mathcal{L}\leftarrow\mathcal{L}+\mathcal{L}_{\mathrm{distill}}$ // Add distillation loss to total loss
25:  end for
26:  $\mathcal{L}_{\mathrm{task}}\leftarrow\mathrm{CELoss}(\hat{\bm{y}}^{s},\text{targets})$ // Compute task loss (e.g., cross-entropy)
27:  $\mathcal{L}_{\mathrm{task}}\leftarrow\text{ranks}[s]\cdot\mathcal{L}_{\mathrm{task}}$ // Weight task loss by student’s rank
28:  $\mathcal{L}\leftarrow\mathcal{L}_{\mathrm{task}}+\beta\cdot\mathcal{L}$ // Combine task loss and distillation loss
29:  return  $\mathcal{L}$ // Return the total loss
Appendix E Classroom Learning Styles Survey

### v2

@@ 4.2 CIFAR-100 Classification with Multiple Mentors
DML 74.83 70.55 70.24 72.86 66.30 74.52
ONE 74.68 70.77 70.43 72.01 66.26 -
SHAKE 75.78 - 71.62 73.85 68.81 76.42
TAKD 75.04 - 70.77 73.67 - -
AEKD 75.68 71.36 71.25 74.75 68.39 76.34
EBKD 74.10 68.24 76.61
DGKD 76.24 - 71.92 74.40 - -
CA-MKD - - - 74.30 69.41 77.94
@@
T: R34, S: R18, 4 P
NOKD 69.75
DML 71.03
ONE 70.55
SHAKE 72.07
KD 70.66
CRD 71.17
AVER 70.63
Ours 71.49
(c) Pose Estimation Results with 4 mentors. We report PCKh for MPII and AP for COCO.
@@
Teacher HRNet-W32-D RTMP-L
@@
Directly aggregating multiple mentors (AVER) often yields modest improvements. ClassroomKD takes this further by ranking and selectively activating mentors, reducing the risk of error accumulation from weaker ones.
Comparison with Specialized Methods.
Techniques like TAKD [5] or DGKD [6] were devised primarily to address large capacity gaps and error propagation, yet ClassroomKD still shows consistently higher accuracy. This underscores the advantage of our fine-grained, per-sample mentor filtering over either purely sequential or random-drop strategies.
Fewer Mentors, Stronger Gains.
Interestingly, we outperform approaches like AEKD [27], which rely on more mentors than we do. This highlights that mentor quality and selective usage are more crucial to final student performance than simply increasing the number of possible teachers.
4.3 ImageNet Classification with Multiple Mentors
To assess scalability, we evaluate ClassroomKD on ImageNet in Table 2(b). ClassroomKD maintains its improvements even when dealing with large-scale data, demonstrating the generality of the rank-based mentor selection. While online SHAKE remains slightly higher, our method still surpasses classic offline KD methods and the naive multi-mentor baseline (AVER).
Partial Online vs. Offline Gap.
Our offline approach does not achieve the same final result as online SHAKE, which benefits from constant inter-model updating. Still, we remain close, suggesting that an adaptive offline framework can approximate or rival online methods without overheads such as co-training multiple large models simultaneously.
4.4 Pose Estimation with Multiple Mentors
Finally, we test ClassroomKD on 2D human pose estimation tasks—both on COCO Keypoints and MPII Human Pose—presented in Table 2(c). Each classroom includes a large high-accuracy teacher (e.g., HRNet-W32-D) plus up to four additional peers. Additional details on our adaptation for pose estimation (e.g., how we compute ranks for heatmap vs. SimCC heads) appear in Appendix B.
Performance Gains.
Unlike classification, pose estimation requires learning structured output (e.g., heatmaps, SimCC x/y logits). Our experiments show that the same rank-based selection and adaptive temperature scaling indeed transfer effectively to these more complex output heads. ClassroomKD consistently achieves better PCKh (MPII) and AP (COCO) than the simplest multi-mentor baseline (AVER). This is mainly attributed to removing guidance from less reliable peers—particularly at the early epochs when the capacity gap is large.
5 Ablation Studies
@@
Table 3: Ablation study to assess the impact of ClassroomKD components.
(a) Role of Multiple Mentors. Single-teacher distillation slightly improves student performance compared to vanilla training. Intermediate mentors (peers) and adaptive distillation further enhances learning.
Student Teacher Peers Adaptive Top-1 Accuracy
@@
5.1 Classroom Size and Composition
@@
11112222333344445555666664646464666666666868686863.3564.7164.7266.5767.0867.53Number of MentorsStudent Top-1S: MBV2, T: EN-B0, 5 Same Peers
@@
Figure 4: Effect of Classroom Size and Composition. We investigate the effect of mentor count, their architectures, and performance differences on learning.
@@
(a) Diversity in mentor architectures. Diverse mentor architectures improve distillation performance compared to a homogeneous setup, even when the parameter count of the diverse mentors is lower. This indicates that architectural diversity provides valuable learning signals.
@@
5.2 Temperature in Mentoring Module
We explore the role of adaptive temperature ( $\tau$) in the Mentoring Module and its impact on bridging the capacity gap between classroom networks. Our approach adjusts the temperature dynamically based on the student’s learning progress, with higher $\tau$ values at the start to accommodate the larger capacity gap, which gradually decreases as the student’s understanding improves (see fig.5). This adaptive strategy allows mentors to effectively “slow down” the teaching process during early stages and accelerate it later, ensuring effective knowledge transfer.
@@ 5.2 Temperature in Mentoring Module
Table 5: Temperature adaption strategy. We compare our temperature adaptation method to DTKD [20] by replacing our mentoring module with their dynamic temperature computation. Our mentoring module outperforms DTKD’s temperature adaption strategy with $\tau=12$ (our default) and $\tau=4$ (tuned for DTKD).
@@
DTKD DTKD 4 69.10 72.05
Ours DTKD 4 64.36 71.18
Ours DTKD 12 68.03 70.02
Ours Ours 12 70.23 72.13
@@
Comparison with DTKD. We compared our approach with DTKD’s dynamic temperature strategy by adding their method to our mentoring module. While DTKD works well with a single teacher (row 1), it is not as effective when used with multiple mentors of different capabilities. This is because DTKD assumes that all mentors predict the correct label and does not fully address the dynamic capacity gap between the teacher and student during the training process. In contrast, our method masks mentor logits with ground-truth labels, and adapts more effectively to evolving capacity gaps, achieving consistently better results across different network architectures.
5.3 Ranking Strategies in KF Module
@@
Table 6: Choice of Ranking Strategy. We compare two ranking methods. Here, we employ different networks for peers. (A) we employ class ranks as $\alpha$, $\beta$ (B) We use discretised probabilities $\lambda$. We observe that using ranks as loss weights improves student network performance compared to probabilities.
Teacher R110 R110 R56 VGG13
Student R20 R32 R20 VGG8
Method B 71.94 74.28 72.56 73.58
Ours(A) 72.06 74.71 72.13 75.29
Teacher VGG13 R32×4 R32×4 R50
Student MBV2 SN-V2 SN-V1 MBV2
Method B 68.52 75.71 75.08 69.78
Ours(A) 70.26 76.74 75.81 70.23
@@ 6 Conclusion
We presented ClassroomKD, a novel knowledge distillation framework that mimics a classroom environment, where a student learns from a diverse set of mentors. By selectively integrating feedback through the Knowledge Filtering (KF) Module and dynamically adjusting teaching strategies with the Mentoring Module, ClassroomKD ensures effective knowledge transfer and mitigates the issues of error accumulation and capacity gap. Our approach significantly improves the student’s performance in classification and pose estimation tasks, consistently outperforming traditional distillation methods.
In large-scale or real-time settings—e.g., mobile deployment for pose estimation—ClassroomKD provides a straightforward mechanism to harness existing high-capacity mentors alongside intermediate peers. As we see in the classification tasks, purely aggregating mentor logits (AVER) or using random dropping is suboptimal. Instead, by dynamically ranking mentors and adjusting teaching temperature per sample and per epoch, we reduce wasted capacity and error buildup.
The main takeaway is that how knowledge is delivered—who is allowed to teach and how swiftly that teaching is introduced—can be just as crucial as which networks are present in the classroom.
Impact.
By fostering more efficient and adaptive students, ClassroomKD paves the way for greener AI solutions with reduced computational costs and energy consumption. Beyond practical applications, this work encourages further research at the intersection of cognitive science and AI, enabling the exploration of more social and educational learning strategies in machine learning. For a discussion of future directions—including applying ClassroomKD to dataset distillation to further reduce memory footprint—please see Appendix E.
Limitations.
While we demonstrated the efficacy of ClassroomKD on image classification and human pose estimation, its application to other domains and more complex tasks, such as object detection and segmentation, presents a promising avenue for future work. Despite the improvements, the framework introduces complexity, especially with respect to the mentor ranking and teaching adjustments, which can require careful tuning. Future work will explore further optimizations and expand the framework’s utility to broader tasks.
Acknowledgement
This study received partial funding from the European Union’s Horizon Europe research and innovation program under grant agreement No. 101135724 (Language Augmentation for Humanverse [LUMINOUS]), addressing Topic HORIZON-CL4-2023-HUMAN-01-21.
Appendix A Analysis and Additional Results
A.1 Per-Class Performance Improvement
Figure 7: Comparison of per-class performance gain over the NOKD baseline. With ClassroomKD (left), the distilled model improves performance on 86 classes. With multi-teacher KD without mentor ranking (right), much fewer classes improve, the absolute improvement is smaller, and the remaining classes experience larger performance degradation (red bars). This highlights the impact of our dynamic strategies in improving performance across different classes.
A.2 Intuition Behind Proposed Ranking Method
Classroom Dynamics. For a given sample $x_{k}$, we can visualize the output probability distribution of a model $m$ by plotting the softmax probability $P_{z_{i}}^{m}$ of its logit $z_{i}$ against the class labels $i$, for all $i\in C$. The models in a classroom can have logit distributions that fall into one of the three cases: (1) Weak classifiers predict the true label $y_{k}$ with low confidence. (2) Strong classifiers predict the true class with high confidence, giving it a "sharper" peak. (3) Wrong classifiers have a peak at the wrong class. This is illustrated in Fig. 8. Using $P_{z_{i}}^{m}$ based ranks will help the student learn to filter out wrong classifiers.
Figure 8: Illustration of the classroom models’ probabilistic distributions The student encounters three types of mentors while learning: 1. weak classifiers predict with low confidence. 2. strong classifiers are highly confident in their prediction. 3. Wrong classifiers predict incorrect labels.
A.3 Dynamic Capacity Gap Visualization
To better understand probabilistic distributions (Fig. 8) of our classroom, we plot the softmax of the logits produced by the student model and mentors at various training steps.
Figure 9: Probability Distributions at Epoch 1. Right subplot is zoomed in at the true class (97).
Figure 10: Probability Distributions at Epoch 121. Right subplot is zoomed in at the true class (4)
Figure 11: Probability Distributions at Epoch 228. Right subplot is zoomed in at the true class (40)
We observe a gradual decrease in the gap between the student and teacher’s probabilities at the true label from epochs 1 through 228.(Fig. 9,  10 and  11)
Appendix B ClassroomKD for 2D Human Pose Estimation
The proposed methodology can be applied to distill knowledge to smaller models in 2D HPE with a few modifications.
B.1 Top-down SimCC-Based Methods
RTMPose architecture, which we use for our experiments on the COCO Kepoints dataset, contains a SimCC [34] head that outputs separate logits of the shape (N, K, D) each in the x and y directions, where N is the batch size, K is the number of joints, and D is the coordinate dimensions. For our purposes, only K is relevant. This output can be seen as two predictions for each of the K joints. Hence, we apply the following three modifications to adapt our approach:
  1. 1.
    The sharpness of model m, $P^{m}$, is calculated using the PCK accuracy metric. These values are further normalized in the classroom to obtain their respective ranks.
  2. 2.
    Once the active mentors are chosen, the $\mathcal{L}_{\mathrm{distill}}$ is processed as the combined distillation loss between the student and mentor along x and y directions.
  3. 3.
    The logits’ shapes are converted to (N*K,-1) before applying the KL-divergence. The sum of distillation losses along the x and y directions is finally divided by the number of joints.
    $\mathcal{L}_{\mathrm{simcc}}(\hat{\bm{y}}^{m},\hat{\bm{y}}^{s};\tau^{m})=\frac{1}{K}(\mathcal{L}_{\mathrm{distill}}(\hat{\bm{y}}_{x}^{m},\hat{\bm{y}}_{x}^{s};\tau^{m})+\mathcal{L}_{\mathrm{distill}}(\hat{\bm{y}}_{y}^{m},\hat{\bm{y}}_{y}^{s};\tau^{m}))$ (15)
B.2 Top-down Heatmap-Based Methods
The LiteHRNet model, which we use for our experiments on the MPII Human Pose dataset, outputs 2D heatmaps of size (N, K, H, W). This is equivalent to the two separate 1D heatmaps in SimCC heads. To apply ClassroomKD in this case, we make the below changes:
  1. 1.
    Similar to the SimCC head, the sharpness of model m, $P^{m}$, is calculated using the PCK metric for the ranking.
  2. 2.
    The KL-divergence between the student and active mentors is calculated between the heatmaps and is then divided by the number of joints.
Appendix C ClassroomKD Algorithm
Algorithm 1 ClassroomKD
0:  Input batch $\mathbf{x}$
0:  Ground truth labels $\mathbf{y}$
0:  Student $s$
0:  Mentors $\mathbb{M}\leftarrow\{t\}\cup\{p_{i}\}_{i=1}^{n}$
0:  $\beta$: weight of distillation loss
0:  $\delta$: weight of standard KD loss with the teacher
3:  $\mathcal{L}\leftarrow 0$ // Initialize total loss
4:  $\mathbb{C}\leftarrow\{s\}\cup\mathbb{M}$
5:  for $m\in\mathbb{C}$ do
6:     $\hat{\bm{y}}^{m}\leftarrow m(\bm{x})$ // Get predictions from model $m$
7:     $\bm{p}^{m}_{\text{gt}}\leftarrow 1/(\text{exp}(\mathrm{CELoss}(\hat{\bm{y}}^{m},\text{targets}))$ // Isolate probabilities assigned to ground truth
[…]
9:     $\text{weights}[m]\leftarrow w^{m}$ // Store weight for model $m$
10:  end for
[…]
14:  for $m\in\mathbb{M}$ do
15:     if $\text{ranks}[m]>\text{ranks}[s]$ then
16:        $\tau^{m}\leftarrow\frac{\text{ranks}[m]-\text{ranks}[s]}{\text{ranks}[m]}$
17:        $\mathcal{L}_{\mathrm{distill}}\leftarrow\mathrm{KL}(\hat{\bm{y}}^{m},\hat{\bm{y}}^{s},\tau^{m})$
18:        $\mathcal{L}_{\mathrm{distill}}\leftarrow\text{ranks}[m]\cdot\mathcal{L}_{\mathrm{distill}}$
19:     else
20:        $\mathcal{L}_{\mathrm{distill}}\leftarrow 0$
21:     end if
22:     $\mathcal{L}\leftarrow\mathcal{L}+\mathcal{L}_{\mathrm{distill}}$ // Add distillation loss to total loss
23:  end for
24:  $\mathcal{L}_{\mathrm{task}}\leftarrow\mathrm{CELoss}(\hat{\bm{y}}^{s},\text{targets})$ // Compute task loss (e.g., cross-entropy)
25:  $\mathcal{L}_{\mathrm{classroom}}\leftarrow\text{ranks}[s]\cdot\mathcal{L}_{\mathrm{task}}+\beta\cdot\mathcal{L}$ // Weight task loss by student’s rank
26:  $\mathcal{L}_{\mathrm{KD}}^{t}\leftarrow\mathcal{L}_{\mathrm{task}}+\mathrm{KL}(\hat{\bm{y}}^{t},\hat{\bm{y}}^{s},\tau^{t}=1)$ // Compute standard student-teacher KD loss
27:  $\mathcal{L}\leftarrow\delta\mathcal{L}_{\mathrm{KD}}^{t}+\beta\cdot\mathcal{L}_{\mathrm{classroom}}$ // Combine KD loss and classroom distillation
28:  return  $\mathcal{L}$ // Return the total loss
Appendix D Training Protocols
@@
Table 7: Mentor configurations used in all our experiments, along with their respective top-1 accuracies and ensemble performance. The size of the mentors, should all the peers be replaced by the teacher ($(n+1)t$), the size of the current mentors ($1tnp$), and the student size are also included.
@@
Model abbreviations: MB: MobileNet, SN: ShuffleNet, R: ResNet, W: WRN, EN: EfficientNet, GN: GoogleNet, RP: RTMPose, HR: HRNet, LHR: LiteHRNet, RG: RegNet
Hardware and Software Configuration. We trained most of our CIFAR-100 experiments on a single V100-16GB GPU. The time required for an experiment ranged between 4 and 4.5 hours on average. We build our code on top of Image Classification SOTA repository [^1] https://github.com/hunto/image˙classification˙sota/ and MMPose, and use pretrained models from these libraries as our mentors.
Appendix E Future Direction: ClassroomKD and Dataset Distillation
@@
Dataset distillation aims to create small, synthetic datasets that enable neural networks to achieve comparable performance to those trained on the original, much larger datasets. This approach reduces computational costs and storage requirements while maintaining model generalization. By optimizing a small set of representative training samples, a distilled dataset $S$ is generated such that a model trained on $S$ performs well on the original dataset $\mathcal{T}$. In our experiments, we use FRePo [35] to create a distilled CIFAR-100 dataset, reducing each class to only 10 samples (Figure 12). Of these, 7 images per class are used for training, while the remaining 3 are used for testing.
@@
Appendix F Classroom Learning Styles Survey